# Mixing Time Bounds for the Gibbs Sampler under Isoperimetry

### v1

@@ 2.2 Assumptions on the target distribution
[…]
@@
Let $q\geq 1$. The distribution $\pi$ is said to satisfy a $L^{q}$-Poincaré inequality ($\mathrm{PI}_{q}$) with constant $\mathtt{p}_{q}>0$ if for any locally Lipschitz function $f:\mathbb{R}^{d}\to\mathbb{R}$,
@@
Let $q\geq 1$. The distribution $\pi$ is said to satisfy a $L^{q}$-log-Sobolev inequality ($\mathrm{LSI}_{q}$) with constant $\mathtt{s}_{q}>0$ if for any locally Lipschitz function $f:\mathbb{R}^{d}\to\mathbb{R}$,
@@
These functional inequalities extend the classical assumption of (strong) log-concavity for $\pi$. Specifically, Bobkov (1999) proved that $\mathrm{PI}_{1}$ is satisfied by all log-concave distributions, and Bakry and Émery (1985) showed that $\mathrm{LSI}_{2}$ holds for all strongly log-concave distributions. As shown in Bobkov and Zegarlinski (2005, Proposition 6.2), $\mathrm{LSI}_{q}$ is also preserved by bounded perturbation à la Holley and Stroock (1987). More generally, Ledoux (1999) proved that $\mathrm{PI}_{2}$ and $\mathrm{LSI}_{2}$ respectively imply sub-exponential and sub-Gaussian tail bounds for the distribution $\pi$. Log-Sobolev inequalites have also been proven for mixture distributions (Chen et al., 2021).
@@
To the best of our knowledge, Proposition 2.4 does not appear explicitly in the literature. For completeness, we include a proof in Appendix A, which may also be of independent interest.
@@
$\left\|\mu-\nu\right\|_{\textsc{TV}}=\sup\left\{\left\lvert\mu(S)-\nu(S)\right\rvert:S\subseteq\mathbb{R}^{d}\text{ measurable}\right\}.$
@@
$\left\|\pi(\cdot\mid x_{-k})-\pi(\cdot\mid y_{-k})\right\|_{\textsc{TV}}\leq M\left\lvert x_{-k}-y_{-k}\right\rvert^{\beta}\qquad\forall k\in[d],~{}\forall x,y\in\mathbb{R}^{d}.$
This type of continuity has been shown to hold for posterior distributions in Bayesian models, under assumptions such as Lipschitz continuity of the likelihood and sufficient integrability of the prior (Dolera and Mainini, 2020, 2023). Moreover, in Subsection 5.1, we will show how to verify that the conditional distributions of $\pi$ are TV-continuous under standard conditions on the potential function $U=-\log\pi$.
@@ 3.1 Mixing times of Markov chains
We begin by recalling key definitions and concepts related to the mixing of Markov chains. Let $\mu$ denote the initial distribution of the Markov chain associated to a kernel $P$. After $k$ steps, the distribution of the chain is given by $\mu P^{k}$, where
@@
It is often assumed that the initial distribution is “not too far” from the stationary distribution, a condition referred to as a warm start.
@@
$\sup\left\{\frac{\mu(S)}{\pi(S)}:S\subseteq\mathbb{R}^{d}\text{ measurable}\right\}\leq\Omega.$
Note that the constant $\Omega$ can grow exponentially with the dimension $d$ even in fairly innocent examples. For instance, this occurs when the initial distribution is a Gaussian centered at the mode of a strongly log-concave target (Dwivedi et al., 2019). Nevertheless, when $\pi$ is strongly log-concave, Altschuler and Chewi (2024) showed that a distribution with a $\mathcal{O}(1)$-warm start can be efficiently obtained using the underdamped Langevin dynamics in $\mathcal{O}(\sqrt{d})$ iterations. We now define the mixing time of a Markov chain as the minimum number of iterations required for the chain to be “close” to its stationary distribution in total variation distance.
@@
$\Phi(P)=\inf\left\{\frac{\int_{S}P(x,S^{c})\pi(\mathrm{d}x)}{\pi(S)}:S\subseteq\mathbb{R}^{d}\text{ measurable},\text{ }\pi(S)\leq\frac{1}{2}\right\}.$
The connection between the mixing time of a Markov chain and its conductance was established by Lovász and Simonovits (1993), who extended the earlier work of Jerrum and Sinclair (1988) to general state spaces. Importantly, this result does not require the transition kernel $P$ to be reversible.
Theorem 3.4 (Lovász and Simonovits, 1993, Corollary 1.5).
Consider a $\pi$-invariant Markov chain with transition kernel $P$ initialized at a $\Omega$-warm start. Then, for all $k\geq 0$,
@@
Consider a $\pi$-invariant Markov chain with transition kernel $P$ initialized at a $\Omega$-warm start. Then, its $\zeta$-mixing time can be bounded by
@@
The problem of upper bounding the mixing time is thus reduced to proving a lower bound on the conductance.
3.2 The spectral gap
Let
$\langle f,g\rangle=\int f(x)g(x)\pi(\mathrm{d}x),$
with corresponding $L^{2}_{\pi}$-norm $\left\|f\right\|_{2}=\sqrt{\langle f,f\rangle}$. Note that a Markov kernel $P$ can be seen as an operator that acts on functions $f\in L^{2}_{\pi}$ via
$Pf(x)=\int f(y)P(x,\mathrm{d}y).$
Let also $I$ denote the identity operator: $If=f$.
Definition 3.6.
Consider a $\pi$-invariant Markov chain with transition kernel $P$. Suppose that $P$ is $\pi$-reversible. The spectral gap of $P$ is defined as
$\lambda_{2}(P)=\inf\left\{\frac{\langle f,(I-P)f\rangle}{\left\|f\right\|_{2}}:f\in L^{2}_{\pi},\text{ }\mathbb{E}[f]=0\right\}.$
When $P$ is reversible, the spectral gap provides a bound on the $L^{2}_{\pi}$-convergence rate of the Markov chain: for any $f\in L^{2}_{\pi}$ with $\mathbb{E}\left[f\right]=0$, we have
$\left\|Pf\right\|_{2}\leq\left(1-\lambda_{2}\right)\left\|f\right\|_{2}.$
See (Levin and Peres, 2017, Exercise 12.4) for a proof. The spectral gap of a reversible Markov kernel $P$ is related to its conductance via the following bounds, also known as Cheeger’s inequalities (Cheeger, 1970).
@@ 3.2 The spectral gap
A lower bound on the conductance thus translates to a lower bound on the spectral gap. Theorem 3.7 applies specifically to the reversible setting and covers the random-scan Gibbs sampler as well as reversibilized forms of the systematic-scan Gibbs sampler, such as the variant with kernel $P_{\mathrm{SS}}^{*}P_{\mathrm{SS}}$ discussed in Subsection 2.1. In this paper, however, we do not directly assume reversibility and instead focus on mixing time bounds as in Corollary 3.5. For non-reversible chains, various notions of spectral gaps have been proposed; see, for instance, Kontoyiannis and Meyn (2012) for definitions based on suitably weighted Banach spaces, and Chatterjee (2023) for finite state spaces. Other approaches (Fill, 1991; Paulin, 2015; Choi, 2020) use reversibilized Markov kernels to derive similar bounds on the convergence rate. However, the sharpness of such bounds in the non-reversible setting remains unclear, and a detailed investigation lies beyond the scope of this work.
@@ 3.3 Bounding the conductance
In order to bound the conductance of a $\pi$-invariant Markov kernel $P$, two main components are required: a three-set isoperimetric inequality for the distribution $\pi$, and a close coupling condition for the kernel $P$. In what follows, we write $\mathrm{Id}:\mathbb{R}\to\mathbb{R}$ for the identity function and use $\sqcup$ to denote the union of disjoint sets.
@@
Let $\Upsilon:(0,\infty)\to(0,\infty)$ and $\Psi:(0,1/2)\to(0,\infty)$ be two non-decreasing functions. Additionaly, assume that $\Psi/\mathrm{Id}$ is non-increasing. The distribution $\pi$ is said to satisfy a three-set isoperimetric inequality with functions $\Upsilon,\Psi$, if for all measurable partitions $S_{1}\sqcup S_{2}\sqcup S_{3}=\mathbb{R}^{d}$,
@@
When $\pi$ is log-concave (or a bounded perturbation thereof), a three-set inequality can be derived by controlling a quantity known as the isoperimetric profile of $\pi$ Milman, 2009b, Theorem 1.2; Andrieu et al., 2024, Section 3. Bounds on the isoperimetric profile have also been established when $\pi$ satisfies $\mathrm{PI}_{1}$ and $\mathrm{LSI}_{1}$ (Rothaus, 1985; Bobkov and Houdré, 1997). In all these cases, $\Upsilon$ is taken to be a multiple of the identity function. However, this approach generally breaks down for non-log-concave $\pi$, as shown in Milman (2009a, Subsection 1.2).
@@
The proof of Proposition 3.10 can be found in Appendix A and builds on several previous works; see Andrieu et al. (2024, Lemma 15) and the survey by Vempala (2005) for further details. Notably, Proposition 3.10 and Corollary 3.11 do not require the Markov kernel $P$ to be reversible. Conventionally, $\delta$ is chosen to scale inversely with the dimension $d$ to ensure that $\varepsilon$ remains of constant order. Under this regime, Corollary 3.11 yields
@@ 4.1 Three-set isoperimetric inequalities via functional inequalities
We now establish three-set isoperimetric inequalities for distributions $\pi$ satisfying an $L^{q}$-Poincaré or $L^{q}$-log-Sobolev inequality for any $q\geq 1$. As noted in Subsection 3.3, traditional methods based on the isoperimetric profile cease to be effective when $\pi$ is not log-concave. To overcome these challenges, we take a more direct approach to derive three-set isoperimetric inequalities for distributions satisfying $\mathrm{PI}_{q}$ or $\mathrm{LSI}_{q}$ with $q\geq 1$. In this setting, $\Upsilon$ is no longer restricted to be proportional to the identity function. In what follows, denote the indicator function of a measurable set $S\subseteq\mathbb{R}^{d}$ as $\mathbbm{1}_{S}$. Additionally, define the $L^{q}_{\pi}$-norm of a function $f$ as
@@
$\Upsilon(t)=\frac{\mathtt{p}_{q}t^{q}}{2^{2q}\left(1+2^{q}\mathtt{p}_{q}t^{q}\right)},\qquad\Psi(t)=t.$
@@
so that $\mathcal{D}(x,S_{1})-\mathcal{D}(y,S_{1})\leq\left\lvert x-y\right\rvert+\xi$. Swapping the roles of $x$ and $y$ and letting $\xi\to 0$, we obtain $\left\lvert\mathcal{D}(x,S_{1})-\mathcal{D}(y,S_{1})\right\rvert\leq\left\lvert x-y\right\rvert$. It follows that the function $f$ is equal to 1 on $S_{1}$, vanishes on $S_{2}$, and is Lipschitz continuous on a subset of $S_{3}$, with the modulus of its gradient equal to $\mathcal{D}(S_{1},S_{2})^{-1}$; elsewhere on $S_{3}$, $f$ vanishes. In particular, $f(x)\in[0,1]$ for $x\in S_{3}$. Applying (2) to $f$, we obtain
$\left\|f-\mathbb{E}[f]\right\|_{q}^{q}=\mathbb{E}\left[\left\lvert f-\mathbb{E}[f]\right\rvert^{q}\right]\leq\frac{1}{\mathtt{p}_{q}}\cdot\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{q}\right]\leq\frac{\pi(S_{3})}{\mathtt{p}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}.$ (4)
@@
$\mathbb{E}\left[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}\right]^{1/q}$ $=\left\|\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\|_{q}$
@@
Using the inequality $\left(a+b\right)^{q}\leq 2^{q}\left(a^{q}+b^{q}\right)$ for $a,b\geq 0$, it follows that
$\mathbb{E}\left[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}\right]\leq 2^{q}\left(2^{q}\left\|f-\mathbbm{1}_{S_{1}}\right\|_{q}^{q}+\left\|f-\mathbb{E}[f]\right\|_{q}^{q}\right).$ (5)
@@
$\left\|f-\mathbbm{1}_{S_{1}}\right\|_{q}^{q}=\mathbb{E}\left[\left\lvert f-\mathbbm{1}_{S_{1}}\right\rvert^{q}\right]\leq\pi(S_{3}).$ (6)
@@
$\mathbb{E}\left[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}\right]\leq 2^{q}\left(2^{q}+\frac{1}{\mathtt{p}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}\right)\pi(S_{3}).$
@@
$\mathbb{E}\left[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}\right]$ $=\left(1-\pi(S_{1})\right)^{q}\pi(S_{1})+\pi(S_{1})^{q}\left(1-\pi(S_{1})\right)$
$=\left(1-\pi(S_{1})\right)\pi(S_{1})\left[\left(1-\pi(S_{1})\right)^{q-1}+\pi(S_{1})^{q-1}\right]$
$\geq\frac{1}{2^{q-1}}\left(1-\pi(S_{1})\right)\pi(S_{1})$
$\geq\frac{1}{2^{q}}\cdot\min\left\{\pi(S_{1}),1-\pi(S_{1})\right\}$
$\geq\frac{1}{2^{q}}\cdot\min\left\{\pi(S_{1}),\pi(S_{2})\right\},$
@@
$\min\left\{\pi(S_{1}),\pi(S_{2})\right\}\leq 2^{2q}\left(2^{q}+\frac{1}{\mathtt{p}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}\right)\pi(S_{3}),$
@@
$\pi(S_{3})\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{2^{2q}\left(\mathtt{p}_{q}^{-1}+2^{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}\right)}\cdot\min\left\{\pi(S_{1}),\pi(S_{2})\right\}.$
@@
$\Upsilon(t)=\frac{\mathtt{p}_{q}t^{q}}{2^{2q}\left(1+2^{q}\mathtt{p}_{q}t^{q}\right)}.$
To derive a three-set isoperimetric inequality corresponding to the $L^{q}$-log-Sobolev inequality, it is useful to analyze the properties of the function $x\mapsto x\cdot\log(1/x)$. The proof of the following lemma is provided in Appendix A.

### v2

@@ 2.2 Assumptions on the target distribution
In the following, let $\left\lvert\cdot\right\rvert$ denote the standard Euclidean norm on $\mathbb{R}^{d}$. Unless stated otherwise, all expectations and variances are taken with respect to the target distribution $\pi$. We say that a function $f:\mathbb{R}^{d}\to\mathbb{R}$ is locally Lipschitz if its restriction to any open ball in $\mathbb{R}^{d}$ is Lipschitz continuous. With these conventions in mind, we now consider two functional inequalities for $\pi$, namely the Poincaré and log-Sobolev inequalities.
@@
Let $q\geq 1$. The distribution $\pi$ is said to satisfy an $L^{q}$-Poincaré inequality ($\mathrm{PI}_{q}$) with constant $\mathtt{p}_{q}>0$ if for any locally Lipschitz function $f:\mathbb{R}^{d}\to\mathbb{R}$,
@@
Let $q\geq 1$. The distribution $\pi$ is said to satisfy an $L^{q}$-log-Sobolev inequality ($\mathrm{LSI}_{q}$) with constant $\mathtt{s}_{q}>0$ if for any locally Lipschitz function $f:\mathbb{R}^{d}\to\mathbb{R}$,
@@
These functional inequalities extend the classical assumption of (strong) log-concavity for $\pi$. Specifically, Bobkov (1999) proved that $\mathrm{PI}_{1}$ is satisfied by all log-concave distributions, and Bakry and Émery (1985) showed that $\mathrm{LSI}_{2}$ holds for all strongly log-concave distributions. As shown in Bobkov and Zegarlinski (2005, Proposition 6.2), $\mathrm{LSI}_{q}$ is also preserved by bounded perturbation à la Holley and Stroock (1987). We recall that a distribution $\nu$ is said to be a bounded perturbation of another distribution $\mu$ if there exist constants $0<c_{1}\leq c_{2}<\infty$ such that the Radon-Nikodym derivative satisfies $c_{1}\leq\mathrm{d}\nu/\mathrm{d}\mu\leq c_{2}$. More generally, Ledoux (1999) proved that $\mathrm{PI}_{2}$ and $\mathrm{LSI}_{2}$ respectively imply sub-exponential and sub-Gaussian tail bounds for the distribution $\pi$. Poincaré inequalities have been established for a broad class of distributions satisfying a Lyapunov condition (Bakry et al., 2008) and log-Sobolev inequalities have also been proven for mixture distributions (Chen et al., 2021).
@@
To the best of our knowledge, Proposition 2.4 does not appear explicitly in the literature. For completeness, we include a proof in Appendix A, which may be of independent interest.
@@
$\left\|\mu-\nu\right\|_{\textsc{TV}}=\sup\left\{\left\lvert\mu(S)-\nu(S)\right\rvert:S\subseteq\mathbb{R}^{d}~\mathrm{measurable}\right\}.$
@@
$\left\|\pi(\cdot\mid x_{-k})-\pi(\cdot\mid y_{-k})\right\|_{\textsc{TV}}\leq M\left\lvert x_{-k}-y_{-k}\right\rvert^{\beta}\qquad\forall k\in[d],\,\forall x,y\in\mathbb{R}^{d}.$
This type of continuity has been shown to hold for posterior distributions in Bayesian models, under assumptions such as Lipschitz continuity of the likelihood and sufficient integrability of the prior (Dolera and Mainini, 2020, 2023). Moreover, in Subsection 5.1, we will show how to verify that the conditional distributions of $\pi$ are TV continuous under standard conditions on the potential function $U=-\log\pi$.
@@ 3 Convergence analysis of Markov chains
3.1 Operator representation and spectral gap
Let
$\left\langle f,g\right\rangle=\int f(x)g(x)\pi(\mathrm{d}x),$
with corresponding $L^{2}_{\pi}$-norm $\left\|f\right\|_{2}=\sqrt{\left\langle f,f\right\rangle}$. Note that a $\pi$-invariant Markov kernel $P$ can be seen as an operator that acts on functions $f\in L^{2}_{\pi}$ via
$Pf(x)=\int f(y)P(x,\mathrm{d}y).$
With this identification, reversibility of $P$ is equivalent to self-adjointness of the associated operator on $L^{2}_{\pi}$. In the reversible setting, the Markov kernel $P$ is said to be positive semi-definite if for all $f\in L^{2}_{\pi}$,
$\left\langle f,Pf\right\rangle\geq 0.$
In particular, the random-scan Gibbs kernel $P_{\mathrm{RS}}$ is positive semi-definite. Indeed, each kernel $P_{k}$ defined in (2) satisfies $P_{k}f(x)=\mathbb{E}[f(X)\mid X_{-k}=x_{-k}]$ and is positive semi-definite since
$\left\langle f,P_{k}f\right\rangle=\mathbb{E}[f(X)\cdot\mathbb{E}[f(X)\mid X_{-k}]]=\mathbb{E}\left[\mathbb{E}\left[f(X)\mid X_{-k}\right]^{2}\right]\geq 0.$
It follows that
$\left\langle f,P_{\mathrm{RS}}f\right\rangle=\frac{1}{d}\sum_{k=1}^{d}\left\langle f,P_{k}f\right\rangle\geq 0.$
A common way to ensure positive semi-definiteness is to work with lazy Markov kernels (see, for example, Lovász and Simonovits, 1993, Subsection 1.b). We say that the kernel $P$ is lazy if
$P(x,\{x\})\geq\frac{1}{2}\qquad\forall x\in\mathbb{R}^{d}.$
Any (reversible or non-reversible) Markov kernel can be made lazy as follows. Let $I$ denote the identity operator, defined by $If=f$.
Definition 3.1.
The lazy version of a Markov kernel $P$ is defined as
$\widetilde{P}(x,\mathrm{d}y)=\frac{1}{2}\delta_{x}(\mathrm{d}y)+\frac{1}{2}P(x,\mathrm{d}y).$
The associated operator is $\widetilde{P}=(I+P)/2$.
In particular, the lazy version of a $\pi$-invariant Markov kernel $P$ is also $\pi$-invariant. An important quantity for analyzing the convergence of reversible Markov chains is the spectral gap, which we now recall.
Definition 3.2.
Consider a $\pi$-invariant Markov chain with transition kernel $P$. Suppose that $P$ is reversible. The spectral gap of $P$ is defined as
$\lambda_{2}(P)=\inf\left\{\frac{\left\langle f,(I-P)f\right\rangle}{\left\|f\right\|^{2}_{2}}:f\in L^{2}_{\pi},\,\mathbb{E}[f]=0,\,f\neq 0\right\}.$
When $P$ is reversible, the spectral gap provides a bound on the $L^{2}_{\pi}$-convergence rate of the Markov chain: for any $f\in L^{2}_{\pi}$ with $\mathbb{E}[f]=0$,
$\left\|Pf\right\|_{2}\leq\left(1-\lambda_{2}(P)\right)\left\|f\right\|_{2}.$
See (Levin and Peres, 2017, Exercise 12.4) for a proof. For non-reversible chains, various notions of spectral gaps have been proposed; see, for instance, Kontoyiannis and Meyn (2012) for definitions based on suitably weighted Banach spaces, and Chatterjee (2025) for finite state spaces. Other approaches (Fill, 1991; Paulin, 2015; Choi, 2020) use reversibilized Markov kernels to derive similar bounds on the convergence rate. However, the sharpness of such bounds in the non-reversible setting remains unclear, and a detailed investigation lies beyond the scope of this work.
@@ 3.2 Mixing times of Markov chains
We begin by recalling key definitions and concepts related to the mixing of Markov chains. Let $\mu$ denote the initial distribution of the Markov chain associated to a kernel $P$. After $k\geq 0$ steps, the distribution of the chain is given by $\mu P^{k}$, where
@@
It is often assumed that the initial distribution is “not too far” from the invariant distribution, a condition referred to as a warm start.
@@
$\sup\left\{\frac{\mu(S)}{\pi(S)}:S\subseteq\mathbb{R}^{d}~\mathrm{measurable},\,\pi(S)>0\right\}\leq\Omega.$
Note that the constant $\Omega$ can grow exponentially with the dimension $d$ even in fairly innocent examples. For instance, this occurs when the initial distribution is a Gaussian centered at the mode of a strongly log-concave target (Dwivedi et al., 2019). Nevertheless, when $\pi$ is strongly log-concave, Altschuler and Chewi (2024a) showed that an initial distribution that is within $\mathcal{O}(1)$ $\chi^{2}$-distance from $\pi$ can be efficiently obtained in $\mathcal{O}(\sqrt{d})$ iterations using the underdamped Langevin dynamics. We note, however, that this notion of proximity for the initial distribution is weaker than the warm start in Definition 3.3. We now define the mixing time of a Markov chain as the minimum number of iterations required for the chain to be “close” to its invariant distribution in total variation distance.
@@
$\Phi(P)=\inf\left\{\frac{\int_{S}P(x,S^{c})\pi(\mathrm{d}x)}{\pi(S)}:S\subseteq\mathbb{R}^{d}~\mathrm{measurable},\,0<\pi(S)\leq\frac{1}{2}\right\}.$
In particular, the conductance of a Markov kernel $P$ and that of its lazy version $\widetilde{P}$ are related by
$\Phi(\widetilde{P})=\frac{1}{2}\cdot\Phi(P).$ (5)
The connection between the mixing time of a Markov chain and its conductance was established by Lovász and Simonovits (1993), who extended the earlier work of Jerrum and Sinclair (1988) to general state spaces.
Theorem 3.6.
Consider a $\pi$-invariant Markov chain with transition kernel $P$ initialized at a $\Omega$-warm start. Suppose that $P$ is either lazy, or both reversible and positive semi-definite. Then, for all $k\geq 0$,
@@
Proof.
If $P$ is lazy, the bound follows from Lovász and Simonovits (1993, Corollary 1.5). If $P$ is reversible and positive semi-definite, the same bound follows from Lee and Zhang (2024, Lemma 5.3), which is a modified version of the former result. ∎
@@
Let $\zeta>0$. Under the conditions of Theorem 3.6, the $\zeta$-mixing time of the Markov chain is bounded by
@@
The problem of upper bounding the mixing time is thus reduced to proving a lower bound on the conductance. The spectral gap of a reversible Markov kernel is also related to its conductance via the following bounds, also known as Cheeger’s inequalities (Cheeger, 1970).
@@
Theorem 3.8 applies specifically to the reversible setting and covers the random-scan Gibbs sampler as well as reversibilized forms of the systematic-scan Gibbs sampler, such as the variant with kernel $P_{\mathrm{SS}}P_{\mathrm{SS}}^{*}$ discussed in Subsection 2.1.
@@ 3.3 Bounding the conductance
In order to bound the conductance of a $\pi$-invariant Markov kernel $P$, two main components are required: an isoperimetric inequality for the distribution $\pi$, and a close coupling condition for the kernel $P$. Following the terminology of Andrieu et al. (2024), we refer to the particular isoperimetric inequalities used here as “three-set” isoperimetric inequalities. In what follows, we write $\mathrm{Id}:\mathbb{R}\to\mathbb{R}$ for the identity function and use $\sqcup$ to denote the union of disjoint sets.
@@
Let $\Upsilon:(0,\infty)\to(0,\infty)$ and $\Psi:(0,1/2)\to(0,\infty)$ be two non-decreasing functions. Additionally, suppose that $\Psi/\mathrm{Id}$ is non-increasing. The distribution $\pi$ is said to satisfy a three-set isoperimetric inequality with functions $\Upsilon,\Psi$, if for all measurable partitions $S_{1}\sqcup S_{2}\sqcup S_{3}=\mathbb{R}^{d}$,
@@
When $\pi$ is log-concave (or a bounded perturbation thereof), a three-set inequality can be derived by controlling a quantity known as the isoperimetric profile of $\pi$ Milman, 2009b, Theorem 1.2; Andrieu et al., 2024, Section 3. Appropriate bounds on the isoperimetric profile are known to be equivalent to $\pi$ satisfying $\mathrm{PI}_{1}$ and, similarly, $\mathrm{LSI}_{1}$ (Rothaus, 1985; Bobkov and Houdré, 1997; Kolesnikov, 2007). In all these cases, $\Upsilon$ is taken to be a multiple of the identity function. However, this approach generally breaks down for non-log-concave $\pi$, as shown in Milman (2009a, Subsection 1.2).
@@
The proof of Proposition 3.11 can be found in Appendix A and builds on several previous works; see Andrieu et al. (2024, Lemma 15) and the survey by Vempala (2005) for further details. Notably, Proposition 3.11 and Corollary 3.12 do not require the Markov kernel $P$ to be reversible. Conventionally, close coupling conditions are established with the parameter $\delta$ scaling inversely with the dimension $d$ and the parameter $\varepsilon$ being of constant order. Under this regime, Corollary 3.12 yields
@@ 4.1 Three-set isoperimetric inequalities via functional inequalities
We now establish three-set isoperimetric inequalities for distributions $\pi$ satisfying an $L^{q}$-Poincaré or $L^{q}$-log-Sobolev inequality for any $q\geq 1$. As noted in Subsection 3.3, traditional methods based on the isoperimetric profile cease to be effective when $\pi$ is not log-concave. To overcome these challenges, we take a more direct approach to derive three-set isoperimetric inequalities for distributions satisfying $\mathrm{PI}_{q}$ or $\mathrm{LSI}_{q}$ with $q\geq 1$. In these settings, $\Upsilon$ is no longer restricted to be proportional to the identity function. In what follows, denote the indicator function of a measurable set $S\subseteq\mathbb{R}^{d}$ as $\mathbbm{1}_{S}$. Additionally, define the $L^{q}_{\pi}$-norm of a function $f$ as
@@
$\Upsilon(t)=\frac{\mathtt{p}_{q}t^{q}}{2^{2q-1}(1+2^{q}\mathtt{p}_{q}t^{q})},\qquad\Psi(t)=t.$
@@
so that $\mathcal{D}(x,S_{1})-\mathcal{D}(y,S_{1})\leq\left\lvert x-y\right\rvert+\xi$. Swapping the roles of $x$ and $y$ and letting $\xi\to 0$, we obtain $\left\lvert\mathcal{D}(x,S_{1})-\mathcal{D}(y,S_{1})\right\rvert\leq\left\lvert x-y\right\rvert$. Hence, the function $f$ is Lipschitz continuous on $\mathbb{R}^{d}$. By Rademacher’s theorem, $f$ is differentiable almost everywhere and satisfies
$\left\lvert\nabla f(x)\right\rvert\leq\frac{1}{\mathcal{D}(S_{1},S_{2})}$
for almost every $x\in\mathbb{R}^{d}$. Moreover, the function $f$ takes the constant value 1 on $S_{1}$, vanishes on $S_{2}$, and satisfies $f(x)\in[0,1]$ for all $x\in S_{3}$. Applying (3) to $f$, we obtain
$\left\|f-\mathbb{E}[f]\right\|_{q}^{q}=\mathbb{E}[\left\lvert f-\mathbb{E}[f]\right\rvert^{q}]\leq\frac{1}{\mathtt{p}_{q}}\cdot\mathbb{E}[\left\lvert\nabla f\right\rvert^{q}]\leq\frac{\pi(S_{3})}{\mathtt{p}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}.$ (6)
@@
$\mathbb{E}[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}]^{1/q}$ $=\left\|\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\|_{q}$
@@
Using the inequality $(a+b)^{q}\leq 2^{q-1}(a^{q}+b^{q})$ for $a,b\geq 0$ and $q\geq 1$, it follows that
$\mathbb{E}[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}]\leq 2^{q-1}\left(2^{q}\left\|f-\mathbbm{1}_{S_{1}}\right\|_{q}^{q}+\left\|f-\mathbb{E}[f]\right\|_{q}^{q}\right).$ (7)
@@
$\left\|f-\mathbbm{1}_{S_{1}}\right\|_{q}^{q}=\mathbb{E}[\left\lvert f-\mathbbm{1}_{S_{1}}\right\rvert^{q}]\leq\pi(S_{3}).$ (8)
@@
$\mathbb{E}[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}]\leq 2^{q-1}\left(2^{q}+\frac{1}{\mathtt{p}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}\right)\pi(S_{3}).$
@@
$\mathbb{E}[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}]$ $=(1-\pi(S_{1}))^{q}\cdot\pi(S_{1})+\pi(S_{1})^{q}(1-\pi(S_{1}))$
$=(1-\pi(S_{1}))\cdot\pi(S_{1})\left[\left(1-\pi(S_{1})\right)^{q-1}+\pi(S_{1})^{q-1}\right]$
$\geq\frac{1}{2^{q-1}}(1-\pi(S_{1}))\cdot\pi(S_{1})$
$\geq\frac{1}{2^{q}}\cdot\min\{\pi(S_{1}),1-\pi(S_{1})\}$
$\geq\frac{1}{2^{q}}\cdot\min\{\pi(S_{1}),\pi(S_{2})\},$
@@
$\min\{\pi(S_{1}),\pi(S_{2})\}\leq 2^{2q-1}\left(2^{q}+\frac{1}{\mathtt{p}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}\right)\pi(S_{3}),$
@@
$\pi(S_{3})\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{2^{2q-1}\left(\mathtt{p}_{q}^{-1}+2^{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}\right)}\cdot\min\{\pi(S_{1}),\pi(S_{2})\}.$
@@
$\Upsilon(t)=\frac{\mathtt{p}_{q}t^{q}}{2^{2q-1}(1+2^{q}\mathtt{p}_{q}t^{q})}.$
To derive a three-set isoperimetric inequality corresponding to the $L^{q}$-log-Sobolev inequality, it is useful to analyze the properties of the function $x\mapsto x\cdot\log(1/x)$. The proof of the following lemma is provided in Appendix A. Throughout, we adopt the convention that $0\cdot\log(0)=0$.